Please explain the below calculation of Pretax cost of Debentures ?


WACC as per book value weights


Book value of Equity & Retained Earnings = Rs.15 crores+ Rs.20 crores =Rs.35 crores

Book Value of preference Shares = Rs. 1 crores

Book Value of Debentures =Rs. 10 crores

Book value of Term loan = Rs.12.5 crores


Cost of Equity = Next year Dividend/Share price + growth rate =3.6/40+0.07 = 0.1600 or 16.00%


Cost of Preference Shares = Dividend/ market price of Preference Share = Rs.100*12%/Rs.75 = 0.1600 or 16.00%


Pretax cost of Debentures(r) is given by (assuming annual coupon)


11.5/r*(1-1/(1+r)^6)+100/(1+r)^6 =80


Solving r =0.1708 or 17.08%



1
Expert's answer
2021-05-30T14:16:25-0400

annual coupon=11.5

maturity=6 years

the nominal price of Debentures=100

discount rate (cost of Debentures)=r

Sold price=80

Formula:

"annual\ coupon/(1+r)+ annual\ coupon/(1+r)^2...annual\ coupon/(1+r)^6+the\ nominal\ price/(1+r)^6= sold\ price"

"annual\ coupon*(\frac{1}{1+r}+...+\frac{1}{(1+r)^6})+the\ nominal\ price/(1+r)^6= sold\ price"

sum of geometric progression: "\frac{\frac{1}{1+r}*(1-\frac{1}{1+r}^6)}{1-\frac{1}{1+r}}=\frac{\frac{1}{1+r}*(1-\frac{1}{1+r}^6)}{\frac{1+r-1}{1+r}}=\frac{1-\frac{1}{1+r}^6}{r}"

"\frac{11.5*(1-\frac{1}{1+r}^6)}{r}+\frac{100}{\frac{1}{1+r}^6}=80"

r=0.1708 or 17.08%